OpenStudy (anonymous):

Solve the differential equation dy/dt = 1/(2y+1)

OpenStudy (anonymous):

(2y+1) dy = 1 dt integrate now

OpenStudy (anonymous):

I just started self studying DIFF EQ yesterday (taking the course once semester starts in 2 weeks) so I'll have a crack at it. (2y+1) dy = 1 dt y^2+ y = t Is that it?

OpenStudy (anonymous):

That was the easy part I need the whole thing y(t) =

OpenStudy (cwrw238):

dont forget the constant of integration

OpenStudy (anonymous):

Here is another one dy/dt = y(1-y) do you know how to do that ? It says the answer is y(t) = ke^t/(ke^t+1)

OpenStudy (anonymous):

ooo right, u(t)=e^t multiply u(t) throughout (e^t * y)'=te^t integrate: ye^t=e^t (t-1) y(t)=t-1+C is that what you got for the 1st question as well?

OpenStudy (anonymous):

No that's not what I got. I figured it out. The answer was (-1 +- sqrt(4t+c))/2

OpenStudy (anonymous):

yah you need to complete the square for first equation

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

Is there any particular reason why you have to complete the square for the 1st?

OpenStudy (anonymous):

cuz you want to solve the function, \(y(t)\) that satisfies the given differential equaiton

OpenStudy (cwrw238):

you have a quadratic equation

OpenStudy (anonymous):

for 2nd question :- \(\large \frac{dy}{dt} = y(1-y)\) \(\large \frac{1}{y(1-y)} dy = dt\) integrate now (left side requires partial fractions)

OpenStudy (cwrw238):

can you split left side into 2 fractions? ita quite straight forward